### v1

Abstract
In this paper we investigate the spatio-temporal dynamics of obesity rates across Italian regions from 2010 to 2022, aiming to identify spatial and temporal trends and assess potential heterogeneities. We implement a Bayesian hierarchical Beta regression model to analyze regional obesity rates, integrating spatial and temporal random effects, alongside gender and various exogenous predictors. The model leverages the Stochastic Search Variable Selection technique to identify significant predictors supported by the data. The analysis reveals both regional heterogeneity and dependence in obesity rates over the study period, emphasizing the importance of considering gender and spatial correlation in explaining its dynamics over time. In fact, the inclusion of structured spatial and temporal random effects captures the complexities of regional variations over time. These random effects, along with gender, emerge as the primary determinants of obesity prevalence across Italian regions, while the role of exogenous covariates is found to be minimal at the regional level. While socioeconomic and lifestyle factors remain fundamental at a micro-level, the findings demonstrate that the integration of spatial and temporal structures is critical for capturing macro-level obesity variations.
keywords:
Bayesian hierarchical model, Gender-specific intercepts, Intrinsic Conditional Auto-Regressive prior, Obesity rates, Stochastic Search Variable Selection
1 Introduction
Obesity is recognized as a major public health problem in Europe, where it has reached alarming levels, posing substantial health, social, and economic burdens on European societies. In 2014, one over six adults were considered obese according to the European Health Interview Survey (2016) [^1] Eurostat news release, 203/2016, 20 October 2016, https://ec.europa.eu/eurostat/documents/2995521/7700898/3-20102016-BP-EN.pdf/c26b037b-d5f3-4c05-89c1-00bf0b98d646. Moreover, projections suggest an increasing average obesity trend (Janssen et al.,, 2020).
Even if the problem is common among countries, such a phenomenon is not homogeneous: overweight and obesity vary substantially between and within countries. For instance, compared to other European countries, Italy shows relatively lower obesity rates for women (OECD,, 2022). According to the Eurostat data (Eurostat,, 2022), the second lowest proportion of obese people in 2019 was recorded in Italy (slightly above 11%), when nearby countries like Spain, France, Germany, and Austria are all above 14 %, and the European mean is around 17%. Despite this, the situation is still problematic and needs appropriate attention and intervention.
Moreover, national averages hide substantial heterogeneity within the countries, both in space and by gender, through time. Focusing on Italy, between northern and southern regions, there is a notable and time consistent difference in overweight and obesity rates. For example, taking the best and worst regions, in the period considered by this study (2010-2022), the female obesity rate (male) is 7.50% (9.40%) in Trentino-Alto Adige and 12.7% (14.4%) in Molise. While considering overweight for the same regions, the percentages are 23.5% (40.3%) for Trentino-Alto Adige and 31% (46%) for Molise. It is well-known that across Italian regions, there are strong structural differences in general culture, social norms, and economic development (for a general discussion, see, for instance, De Blasio and Nuzzo,, 2009). The diversity observed among Italian regions may influence obesity and related phenomena (Gallus et al.,, 2013). For this reason, proposing an analysis that takes into account heterogeneity is essential (Brunello and Labartino,, 2014). In a related direction, flexible models - such as varying coefficient models - have also been proposed to explore time-varying and spatially-varying effects on public health outcomes, including obesity, particularly using large-scale survey data such as the Behavioral Risk Factor Surveillance System (Assaf et al.,, 2015). Furthermore, gender aspects should also be considered, given potential physical and genetic differences (Varì,, 2016) as well as cultural ones (Barra and Barone,, 2021).
In this paper, we analyse longitudinal areal data consisting of yearly regional obesity rates from 2010 to 2022. To model these rates, expressed as proportions bounded between 0 and 1, we employ a Beta regression model within a Bayesian hierarchical framework. This approach provides a flexible structure for modelling rate data, enabling the inclusion of covariates as well as unobserved spatial and temporal heterogeneity through structured random effects. The Beta regression framework has been extensively studied over the past two decades. For instance, Simas et al., (2010) extend Beta regression by modelling both the mean parameter and the precision parameter as linear combinations of covariates, linked via appropriate functions. Similarly, Da Silva and De Oliveira Lima, (2017) incorporate Beta regression into a geographically weighted regression framework, adapting the model to account for spatial variation. Furthermore, Figueroa-Zúñiga et al., (2013) extend Beta regression by embedding it within a Bayesian hierarchical framework, enabling the inclusion of random effects. Our Beta regression model follows the Bayesian hierarchical framework proposed by Figueroa-Zúñiga et al., (2013), but introduces key differences in the specification of random effects and in the approach used for variable selection. These modifications are designed to address the specific challenges of modelling spatio-temporal structures in areal data and to facilitate effective selection from a large set of covariates. Bayesian Beta regression models with spatio-temporal random effects have been explored in the literature, such as in the work of Lagos-Alvarez et al., (2017), which focuses on geostatistical data rather than areal units. However, to the best of our knowledge, no existing Bayesian Beta regression model for areal data simultaneously incorporates spatio-temporal random effects, group-specific intercepts, and Bayesian variable selection. We mention that other approaches to deal with rate data have been studied. For example, compositional data analysis has been proposed to jointly forecast multiple weight categories (e.g., underweight, overweight, obese) using log-ratio transformations (Mills,, 2008).
To account for spatial and temporal correlation, we incorporate an additive formulation for the random effects in the linear predictor of the Beta regression model. Specifically, we assume an Intrinsic Conditional Auto-Regressive (ICAR) prior for the spatially structured random effect, exploiting its ability to capture structured spatial dependence, which is particularly relevant for regional-level data where geographic proximity strongly influences observed outcomes (Banerjee et al.,, 2014; Haining and Li,, 2020). The temporal random effect is modelled as a first-order auto-regressive process (AR(1)). Furthermore, in order to capture potential group differences across gender, we include gender-specific intercept terms in the model. This allows to assess the influence of gender alongside spatial and temporal effects. In this way, our approach jointly considers space, time, and gender influences on regional obesity rates, all of which have been shown to be highly relevant (Grimaccia and Rota,, 2025). The additive combination of the spatial and temporal random effect offers a parsimonious and interpretable framework, making it a suitable modelling strategy (Banerjee et al.,, 2014). Although more structured interactions between spatial and temporal components could be incorporated (for a review, see Knorr-Held,, 2000), we follow the approach of Waller et al., (1997) and Berloco et al., (2023), arguing that prioritizing simplicity and interpretability over added flexible and complexity aligns with the primary objective of this study: to effectively model regional-level patterns in obesity rates while focusing on structured spatio-temporal dynamics.
With the objective of finding a parsimonious representation of the observed data using a statistical model that is also capable of accurate prediction, we adopt the Stochastic Search Variable Selection (SSVS) approach (George and McCulloch,, 1993, George and McCulloch,, 1997). This method is a Bayesian regularization technique that has proven effective in addressing both prediction and variable selection problems (see Rockova et al.,, 2012, for a detailed discussion). SSVS is particularly attractive because it is well-suited for high-dimensional modelling, provides valid standard errors, and enables the simultaneous estimation of regression coefficients and complexity parameters through computationally efficient Markov chain Monte Carlo (MCMC) techniques.
The paper is structured as follows. In Section 2, we detail the construction of the dataset, which was assembled from multiple official sources. In Section 3 we outline the Bayesian hierarchical spatio-temporal Beta regression model (Section 3.1), providing details about prior assumptions (Section 3.2) for the structured spatial and temporal random effects, alongside the gender-specific intercepts. Section 3.3 and 3.4 discuss the SSVS prior specification for the fixed effects and some computation details, respectively. We then present the results of our analysis in Section 4 and some final considerations in Section 5. Additional material is provided in the appendices: Appendix 10 contains the table describing the covariates used in the analysis; Appendix 11 reports additional posterior summary statistics not included in the main text for brevity; and Appendix 12 presents the results of the comparison of some variants of the main model.
2 Data
The dataset analysed in this study comprises spatio-temporal areal data, which can be conceptualized as a set of time series associated with each spatial unit (i.e., region). The dataset is organized as a longitudinal dataset, collecting information on the 20 Italian regions from 2010 to 2022 (13 years). The Italian regions are divided, according to the NUTS 1 classification [^2] NUTS classification: https://ec.europa.eu/eurostat/web/nuts into 5 different geographical zones: North West (NO), North East (NE), Centre (C), South (S) and Islands (SI). In Table 1 we report the regions and the corresponding geographical zone.
Table 1: Italian regions grouped by geographical zone. Source: Geographical classification of Italian regions according to the Italian National Institute of Statistics (ISTAT).
Zone Regions
North West (NO) Piemonte, Valle d’Aosta, Lombardia, Liguria, Emilia-Romagna
North East (NE) Trentino-Alto Adige, Veneto, Friuli-Venezia Giulia
Centre (C) Toscana, Umbria, Marche, Lazio
South (S) Abruzzo, Molise, Campania, Puglia, Basilicata, Calabria
Islands (SI) Sicilia, Sardegna
[…]
The dataset contains aggregated data at the regional level (the NUTS 2 level) with variables belonging to four groups: socio-demography, mortality, lifestyle, and general health-related information (see Table 7 in Appendix 10). We assembled the dataset by extracting information from the Health For All platform (HFA [^3] HFA: https://www.istat.it/it/archivio/14562), the Benessere Equo e Sostenibile (BES [^4] BES: https://www.istat.it/en/well-being-and-sustainability/the-measurement-of-well-being/bes-report) report, the Eurostat databases [^5] Eurostat: https://ec.europa.eu/eurostat/data/database and the archive of the Ministry of Economy and Finance (MEF [^6] MEF archive: www.rgs.mef.gov.it/VERSIONE-I/archivio).
The dependent variable is the obesity rate measured yearly for each region in the time window considered separately by gender (source: HFA). The definition of obesity follows the WHO classification [^7] WHO Obesity definition: https://www.who.int/news-room/fact-sheets/detail/obesity-and-overweight, i.e., having a Body Mass Index (BMI) greater than or equal to 30 defines an obese individual.
For the mortality rate variables in Group 2 of Table 7, which exhibited high multicollinearity and strong linear relationships, we applied a principal component analysis (PCA). These variables represent cause-specific mortality rates, calculated as the number of deaths due to a given cause per 10,000 individuals within the same age and gender group. Among the nine principal components generated, we retained the first two, which together explained more than 70% of the total variance. The first principal component (PC1) primarily captures diseases related to cardiovascular issues and diabetes, while the second (PC2) reflects factors more associated with stress and mental disorders.
3 Methods
3.1 Beta regression model
Let $Y_{its}$ denote the obesity rate in region $i$ $(i=1,\dots,20)$, year $t$ $(t=1,\dots,13)$ and gender $s$ $(s=1\text{ for females and}$ $s=2\text{ for males})$. We specify the following Beta regression model (Ferrari and Cribari-Neto,, 2004):
$Y_{its}\mid\mu_{its},\phi\stackrel{{\scriptstyle ind}}{{\sim}}Beta(\mu_{its}\phi,\ \ (1-\mu_{its})\phi)$
$\ln\left(\frac{\mu_{its}}{1-\mu_{its}}\right)=\eta_{its}$
with the linear predictor $\eta_{its}$ defined as:
$\eta_{its}=\beta_{0}+\bm{x}_{its}^{T}\ \bm{\beta}+\psi_{i}+\alpha_{t}+d_{s}\gamma.$
With this formulation we have that $\mu_{its}=E(Y_{its})\in(0,1)$ and $\text{Var}(Y_{its})=\left(\mu_{its}(1-\mu_{its})\right)/(1+\phi)$. It follows that $\mu_{its}$ is the mean obesity rate for region $i$ at year $t$ and for gender $s$, and the parameter $\phi$ can be interpreted as a constant precision parameter, i.e., given $\mu_{its}$, the larger the value of $\phi$, the smaller the variance of $Y_{its}$. The structured space random effect is represented by $\bm{\psi}=(\psi_{1},\psi_{2},\ldots,\psi_{20})^{T}$, while the structured time random effect by $\bm{\alpha}\ =(\alpha_{1},\alpha_{2},\ldots,\alpha_{13})^{T}$. The global intercept term is $\beta_{0}$, while the gender-specific intercepts are introduced through the dummy variable $d_{s}$ which takes the value 0 for females and 1 for males. Following Berloco et al., (2023), we adopt an additive specification for the spatial and temporal effects. More details about the associated prior specification are provided in Section 3.2. The fixed effects are introduced in the model through the vector of $p=32$ exogenous covariates (see Table 7 in Appendix 10), denoted by $\bm{x}_{its}=(x_{its1},x_{its2},\ldots,x_{itsp})^{T}$, and the corresponding vector of coefficients $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{p})^{T}$.
The Beta parametrization adopted in our model offers distinct advantages. First, by expressing the distribution in terms of its mean $\mu_{its}$ and precision $\phi$, the model facilitates a more intuitive interpretation. Second, this formulation simplifies the specification of prior distributions within the Bayesian framework. Prior knowledge can be naturally incorporated through $\mu_{its}$ via the linear predictor. Simultaneously, the precision parameter $\phi$ can be assigned a weakly informative prior, allowing the model to flexibly adapt to the data’s inherent variability (Ferrari and Cribari-Neto,, 2004).
3.2 Prior Specification
To address the structured spatial random effect within a Bayesian hierarchical framework, we assume an ICAR prior for the random vector $\bm{\psi}$. Specifically, the spatially structured random effect is modelled as follows:
$\bm{\psi}\mid\tau_{\psi}\sim\mathcal{N}\big{(}\bm{0},[\tau_{\psi}(\bm{D}-\bm{W})]^{-1}\big{)},$
where $\bm{W}$ is a binary symmetric $20\times 20$ matrix known as the “weights matrix” or “adjacency matrix” (see, e.g., Banerjee et al.,, 2014; Haining and Li,, 2020). The entries of $\bm{W}$ are defined as follows: $w_{ij}=1$ if regions $i$ and $j$ are neighbours, and $w_{ij}=0$ otherwise. Diagonal elements are set to zero since a region is not considered a neighbour of itself. The diagonal matrix $\bm{D}$ is defined as $\bm{D}=\textbf{Diag}(\bm{W}\cdot\bm{1})$, where Diag represents the diagonal matrix operator, $\bm{1}$ denotes a column vector of ones and $\cdot$ denotes the dot product. The diagonal entries $D_{ii}$ represent the number of neighbours for each region $i$. The scalar $\tau_{\psi}$ is a precision parameter that scales the degree of spatial smoothing in the model: higher values of $\tau_{\psi}$ lead to stronger spatial smoothing, while lower values allow for greater heterogeneity across regions. The full-conditional distribution of the $i$-th spatial random effects $\psi_{i}$ is given by:
$\psi_{i}\mid\psi_{j\neq i},\tau_{\psi}\sim\mathcal{N}\bigg{(}\sum_{j\sim i}\frac{\psi_{j}}{D_{ii}},\frac{1}{D_{ii}\tau_{\psi}}\bigg{)},$ (1)
where $j\sim i$ indicates that regions $i$ and $j$ are neighbours. As shown in Equation (1), the conditional mean of $\psi_{i}$ is the average of the values of its neighbours, while its variance decreases as the number of neighbours increases. This formulation implies complete spatial dependence, which means that the value of each region is entirely determined by the values of its neighbours (Rue and Held,, 2005).
For what concerns the structured time random effects, following Knorr-Held, (2000) and Berloco et al., (2023), we model the temporal random vector $\bm{\alpha}$ using a first-order auto-regressive (AR(1)) process with the following structure:
$\alpha_{1}\mid\tau_{\alpha}\sim\mathcal{N}(0,\tau_{\alpha}^{-1}),$
$\alpha_{t}\mid\alpha_{t-1},\rho,\tau_{\alpha}\sim\mathcal{N}(\rho\alpha_{t-1},\tau_{\alpha}^{-1}),\quad t=2,\ldots,13,$
where $\tau_{\alpha}$ is the precision parameter and $\rho$ is the auto-regressive coefficient. We assign a uniform prior to $\rho$:
$\rho\sim Beta(1,1),$
such that $\rho$ is restricted to the interval $[0,1]$, reflecting a prior belief in positive autocorrelation among the temporal random effects while preventing explosive dynamics in the AR(1) process.
The gender-specific intercepts are modelled in an additive way. First, we recall that the individual $\psi_{i}$’s cannot be uniquely determined. This issue is resolved by imposing a sum to zero constraint $\sum_{i=1}^{n}\psi_{i}=0$ (Thomas et al.,, 2014) and by modelling the global intercept term $\beta_{0}$ as suggested by Besag and Kooperberg, (1995):
$p(\beta_{0})\propto 1,\quad\text{for}\quad\beta_{0}\in(-\infty,+\infty),$
indicating a flat, improper prior for $\beta_{0}$ over the entire real line. A gender-specific shift is introduced through the binary indicator variable $d_{s}$, where $d_{s}=0$ for females and $d_{s}=1$ for males. The gender effect coefficient $\gamma$ allows the definition of gender-specific intercepts as
$\xi_{s}=\beta_{0}+d_{s}\gamma,$
so that $\xi_{1}=\beta_{0}$ for females and $\xi_{2}=\beta_{0}+\gamma$ for males. We assign to $\gamma$ a weakly informative normal prior:
$\gamma\sim\mathcal{N}(0,\sigma^{2}_{\gamma}=100).$
For the structured spatial and temporal random effects, we assign a Gamma prior to the respective precision parameters. Following the recommendation of Kelsall and Wakefield, (1999), we use a Gamma with shape and rate parameters $\alpha_{\tau}=0.5$ and $\beta_{\tau}=0.0005$ for the spatial precision $\tau_{\psi}$, as it reflects a conservative and weakly informative belief about spatial dependence (Thomas et al.,, 2014). Although originally proposed for spatial effects, we extended this prior to the temporal precision $\tau_{\alpha}$, acknowledging the presence of temporal correlation in our setting, favoring shrinkage while preserving flexibility to capture stronger spatial or temporal structure when supported by the data:
$\tau_{\psi},\tau_{\alpha}$ $\sim\text{Gamma}(\alpha_{\tau},\beta_{\tau}).$
Finally, we assign a prior to the Beta precision parameter $\phi$ following the approach proposed by Figueroa-Zúñiga et al., (2013), which extends the prior structure suggested by Gelman, (2006) specifically for the Beta distribution, by assuming:
$\phi=(aB)^{2},\quad B\sim\text{Beta}(1+\epsilon,1+\epsilon),$
where $a$ is a large constant (e.g., 50) and $\epsilon$ is a small constant (e.g., 0.1).
To enable variable selection, a prior on the covariate coefficients $\bm{\beta}$ is specified following the SSVS framework, as detailed in Section 3.3.
Table 2: Summary of the model parameters.
Parameter Description
$\xi_{s}$ Gender-specific intercept term ($s=1,2$)
$\psi_{i}$ Structured spatial random effect ($i=1,\ldots,20$)
$\tau_{\psi}$ Precision parameter for the spatial random effect
$\alpha_{t}$ Structured temporal random effect ($t=1,\ldots,13$)
$\tau_{\alpha}$ Precision parameter for the temporal random effect
$\rho$ AR(1) coefficient for the temporal random effect
$\phi$ Beta precision parameter
$\beta_{k}$ Coefficient for the exogenous variables ($k=1,\ldots,p$)
$\sigma_{k}$, $\omega_{k}$, $\theta_{k}$ Parameters from the SSVS prior
[…]
3.3 Stochastic Search Variable Selection Prior
To model the fixed effects coefficients $\beta_{k},k=1,\dots,p$, we adopt a SSVS prior specification (George and McCulloch,, 1993). This approach combines a spike-and-slab mixture prior to probabilistically determine covariate inclusion. In a nutshell, SSVS enables each regressor to be significantly associated with the response (i.e., different from zero) or shrunk toward zero. This is achieved by assigning a prior that assumes independence between regressors, with marginals defined as a mixture of two Normal distributions, both centered at zero. The spike component corresponds to negligible or irrelevant effects, while the slab component allows for greater variance and captures substantial or practically meaningful associations. Specifically, we assign the following hierarchical prior structure:
$\beta_{k}\mid\sigma^{2}_{k}$ $\sim\mathcal{N}(0,\sigma_{k}^{2}),$
$\sigma_{k}^{2}\mid\omega_{k}$ $=(1-\omega_{k})\tau^{2}+\omega_{k}\,c^{2}\tau^{2},$
$\omega_{k}\mid\theta_{k}$ $\sim\text{Bernoulli}(\theta_{k}),$
$\theta_{k}$ $\sim\text{Uniform}(0,1).$
In this formulation, $\omega_{k}$ is a binary latent variable indicating whether the $k$-th covariate is “practically significant” ($\omega_{k}=1$) or not. The prior probability of inclusion of the $k$-th covariate is denoted by $\theta_{k}$, to which we assign a Uniform hyper-prior, avoiding imposing strong prior assumptions about the relevance of individual predictors. The parameters $\tau^{2}$ and $c^{2}\tau^{2}$ represent the variances of the spike and slab components, respectively. The intersection points of the spike and slab densities define a threshold $\zeta$ for practical significance: $\zeta=\pm\tau\sqrt{2c^{2}\log(c)/(c^{2}-1)},$ such that coefficients falling within the interval $\zeta,+\zeta]$ are interpreted as practically negligible. Therefore, in accordance with Rockova et al., (2012), we fix to constant values $\tau$ and $c$ to control the separation between negligible and relevant effects (see Section 3.4). Variable selection is then achieved in practice by either identifying the model with the highest posterior probability, including or excluding covariates based on their marginal posterior inclusion probabilities, or evaluating the extent to which the marginal posterior distributions of the regression coefficients are shrunk toward zero (for further details see Rockova et al.,, 2012).
3.4 Computational Details
A summary of all the model parameters is reported in Table 2. To conduct Bayesian inference, we implemented the proposed spatio-temporal hierarchical model using the Nimble package (de Valpine et al.,, 2017, 2024), which enables efficient specification and sampling of complex models. The final model configuration was determined after extensive empirical testing to ensure posterior convergence and computational efficiency. To improve convergence of the MCMC algorithm, we normalized the variables using min-max scaling and standardization, as appropriate. For what concerns the SSVS specification, we set the slab scaling factor to $c=4000$, and the intersection term to $\zeta=0.001$, reflecting the threshold of practical significance we determined through iterative exploration and consideration of the magnitude of the coefficients. We derived the standard deviation of the spike using the relationship: $\tau={\zeta}/{\sqrt{{2\log(c)c^{2}}/(c^{2}-1)}},$ which yielded $\tau=0.00025$, a value corresponding to a very small standard deviation for the spike around 0. The slab variance was then calculated as $\tau^{2}c^{2}$, resulting in $\tau^{2}c^{2}\approx 1$. This configuration aligns the slab variance with the suggested scale of variation for coefficients under the logit link function, as suggested by McElreath, (2016).
We performed parameter estimation via the MCMC algorithm in Nimble, with 150,000 iterations across two chains, a burn-in period of 50,000 iterations and a thinning interval of 10, yielding a 10,000 posterior sample size.
4 Results
Table 3 presents the posterior summary statistics for the gender-specific intercepts, which are defined as $\xi_{1}=\beta_{0}$ for females and $\xi_{2}=\beta_{0}+\gamma$ for males (see Section 3.2). The mean of the intercept for females is 1.09-1.09, indicating a lower baseline effect compared to males, whose corresponding mean is $0.33$. The standard deviation is slightly higher for males ($0.18$) than for females ($0.16$). This difference in variability is further reflected in the coefficients of variation (CV), suggesting a slightly higher sampling variability for the male intercept. Both distributions are approximately symmetric. The kurtosis values are close to that of a normal distribution for males ($3.03$), whereas the female-specific distribution exhibits slightly heavier tails ($3.48$).
Table 3: Posterior summary statistics for the gender-specific intercepts $\xi_{1}$ and $\xi_{2}$.
Parameter (Gender) Mean SD CV Skewness Kurtosis
$\xi_{1}$ (Females) -1.09 0.16 -0.15 -0.22 3.48
$\xi_{2}$ (Males) 0.33 0.18 0.53 -0.11 3.03
[…]
Figure 1: Posterior distributions of the gender-specific intercepts: $\xi_{1}$ for females (in pink) and $\xi_{2}$ for males (in light blue).
Figure 1 depicts the posterior density estimates of the gender-specific intercepts. The clear separation between the two distributions emphasizes the significant role gender plays in modeling obesity. This result highlights the substantial contribution of gender-specific effects in explaining differences in obesity prevalence.
Figure 2: Posterior density estimation of the space random effects $\psi_{i}$, coloured by geographical zone (NO=North West, NE=North East, C=Centre, S=South, SI=Islands, see Table 1).
Figure 3: Posterior mean values of the spatial random effects $\psi_{i}$ for each Italian region.
The structured spatial random effects can be visualized through the posterior densities plotted in Figure 2. The plot shows the posterior densities for all regions, highlighting the contribution of regional random effects to the overall obesity estimates according to the model. Overall, the ICAR prior effectively induces a spatial structure that captures regional dependencies while allowing for evident heterogeneity, as reflected in the varying magnitudes of $\psi_{i}$ across macro areas and individual regions. To complement the density plots, Figure 3 provides a map representation of the posterior mean values of the spatial random effects, facilitating the interpretation of their spatial distribution. While spatial proximity generally corresponds to similar values of the spatial random effect, notable exceptions suggest that additional determinants beyond geography influence regional obesity rates. As shown by the posterior summary statistics in Table 8 of Appendix 11, regions with negative posterior mean values, such as Piemonte (-0.61), Liguria (-0.62), Trentino-Alto Adige (-0.49), and Sardegna (-0.71, the lowest), are associated with lower obesity rates. In contrast, with the exception of Calabria, all southern regions, particularly Molise (0.89, the highest), as well as Umbria (0.29), exhibit strong positive spatial effects. The uncertainty in spatial effects varies across regions, as reflected in the standard deviations of the posterior distributions. Trentino-Alto Adige and Calabria (SD = 0.12, the highest), Campania, and Sicilia exhibit greater variability, whereas Veneto, Friuli-Venezia Giulia, Umbria, Marche, Lazio and Abruzzo show more spiked posterior estimates, with Abruzzo having the lowest SD (0.06). Additionally, the distributions reveal a skewness pattern: northern regions such as Piemonte, Lombardia, Liguria, and Lazio exhibit right-skewness, while southern regions generally show left-skewness. Notably, all the regions have kurtosis values exceeding 3, suggesting pronounced posterior tails, with Puglia being the region with the heaviest tails. To summarize, while spatial correlation in obesity rates is evident, the variation across regions is not as smooth as one might expect from a prior ICAR specification. Some regional posterior densities deviate strongly from those of neighbouring regions, suggesting localized spatial heterogeneity that is not entirely captured by the prior smoothing assumptions. As regards the posterior summary statistics for the spatial precision parameter $\tau_{\psi}$, we refer the reader to Appendix 11.
Figure 4: Posterior density estimation of the time random effects $\alpha_{t}$, coloured by year.
Posterior density estimates of the time-specific random effects are shown in Figure 4. A clear temporal pattern emerges: during the first half of the observational window (2010–2015), the random effects remain relatively stable and are centered around values slightly below zero. In contrast, the second half of the period exhibits a marked upward shift in the magnitude of the effects. Between 2010 and 2015, the posterior probability that the temporal effects exceed zero remains relatively stable, fluctuating around 0.3, with the exception of 2014 which shows a modest increase. Starting in 2016, these probabilities rise significantly: both 2016 and 2017 reach approximately 0.76, and from 2018 onward, the vast majority (if not all) of posterior samples are positive. These trends are summarized in Table 4.
Table 4: Posterior probability that the temporal effect $\alpha_{t}$ exceeds zero (PP0).
$t$ 1 2 3 4 5 6 7 8 9 10 11 12 13
Year 2010 2011 2012 2013 2014 2015 2016 2017 2018 2019 2020 2021 2022
PP0 0.28 0.25 0.25 0.38 0.61 0.32 0.76 0.76 0.96 0.98 1.00 1.00 0.99
[…]
As confirmed by the posterior summary statistics shown in Table 9 of Appendix 11, the mean values of the $\alpha_{t}$ parameters range from negative values, such as $\alpha_{2}$ and $\alpha_{3}$, which have the lower mean (-0.08), to positive values, with $\alpha_{12}$ exhibiting the highest one (0.41). These differences highlight the varying influence of the time random effects across different time points. While some periods exhibit negative random effects, indicating a lower-than-expected influence on the obesity rates, others demonstrate positive effects, which may indicate periods of higher-than-expected influence. The standard deviations of the $\alpha_{t}$ parameters are relatively stable, ranging from 0.12 to 0.15. This consistency suggests that while the magnitude of time random effects may differ across time points, the variability within each time point is relatively homogeneous. With few exceptions, the $\alpha_{t}$ coefficients show a positive skewness, consistent in the second half of the period, indicating again the increase in the temporal effect. All of the $\alpha_{t}$ parameters exhibit kurtosis values greater than 3, interestingly showing the heavy tails of all the posterior distributions of the time random effect.
Table 5: Posterior summary statistics for the auto-regressive coefficient $\rho$.
Parameter Mean SD CV Skewness Kurtosis
$\rho$ 0.85 0.14 0.16 -1.48 5.62
[…]
We briefly mention that the auto-regressive parameter $\rho$ show a negatively skewed posterior distribution, with a posterior mean of 0.85, suggesting quite strong temporal dependence from year to year (see Table 5). As regards the posterior summary statistics for the temporal precision parameter $\tau_{\alpha}$, we refer the reader to Table 10 in Appendix 11.
Table 6: Posterior summary statistics for the Beta precision parameter $\phi$.
Parameter Mean SD CV Skewness Kurtosis
$\phi$ 37.67 2.39 0.06 0.14 3.00
[…]
The posterior summary statistics of the Beta precision parameter $\phi$ reported in Table 6 are not straightforward to interpret, primarily because they depend on the scale of the data. The parameter becomes more interesting when compared between models: a model with a higher posterior precision parameter suggests a better ability to explain the variability of the Beta-distributed response variable, due to the relation that links the variance to the precision $\phi$ (see Section 3). We just highlight that the Bayesian updating process was very pronounced, as the prior distribution, which spans a wide range (between 0 and $2500$), is effectively updated to a more concentrated range around the mean of 37.67. The reduction in uncertainty is very high, as reflected by the low coefficient of variation equal to 0.06. Finally, we observe only a mild positive skewness and a kurtosis value nearly identical to that of a Normal distribution.
Figure 5: Posterior probability evaluation of the inclusion parameters $\omega_{k}$ for all the covariates (for the description of the covariates see Table 7 in Appendix 10).
As discussed in Section 3.3, we adopt the SSVS approach to perform inference on the fixed-effect coefficients $\beta_{k}$. Consequently, the evaluation of their posterior distribution is intrinsically linked to the variable selection induced by the spike and slab prior. The SSVS hierarchical prior specification can make inference on the $\beta_{k}$ coefficients challenging, particularly when their posterior distributions are bi-modal, with samples alternating between values near zero and substantially different from zero. This complicates the inferential process. However, we point out that the primary objective of this work is to investigate the spatio-temporal dynamics of obesity rates across Italian regions, rather than exploring or assessing causal relationships between the covariates and the response variable. The inclusion of these covariates is primarily to control for any additional effect that may arise along with the random effects, which are of greater interest in this study. Thus we do not conduct detailed posterior evaluation of the $\beta_{k}$ coefficients. Nevertheless, given the valuable insights offered by the SSVS approach, we perform a brief analysis of the inclusion probabilities to assess which covariates are associated with obesity, after accounting for spatial and temporal random effects, as well as gender. Specifically, the $\omega_{k}$ parameter plays a central role in determining whether the $k$-th covariate should be included in the model, based on the probability parameter $\theta_{k}$ (see Section 3.3). By counting the proportion of MCMC iterations in which $\omega_{k}=1$, we obtain an empirical inclusion probability for each covariate. Figure 5 presents these posterior inclusion probabilities. The only covariates that are consistently included in the model are life expectancy, life expectancy in good health and overweight. All the remaining covariates exhibit very low posterior inclusion probabilities. This suggests that, once the random effects and gender-specific intercepts are incorporated, only three covariates need to be included according to our model specification. The other covariates do not prove to be related with obesity rates and only add unnecessary complexity to the model.
In order to assess the performance of our modelling framework, we considered three alternative model specifications alongside the main model presented in Section 3. These variants differ in two aspects: (i) the treatment of the gender effect, modelled either as a fixed or a random effect; and (ii) the prior distribution on the auto-regressive coefficient $\rho$, specified either as a Beta(1,1)(1,1) distribution (constraining $\rho$ to the unit interval) or as a standard Normal distribution (allowing unrestricted support). This results in four model configurations, named M1, M2, M3 and M4, as detailed in Table 11 in Appendix 12. Note that M1 (i.e., gender as fixed effect and Beta prior for $\rho$) corresponds to our main model described in Section 3. To compare the predictive performance of these four models, we performed a temporal training-test split, using data from 2022 ($t=13$) as the test set and the previous years ($t=1,\dots,12$) for training. Models were assessed using standard predictive metrics. The results indicate that M1 outperforms the competing models in terms of predictive accuracy (see Table 12 in Appendix 12). The latter is likely due to the temporal dynamics allowed by the Normal prior, which can lead to explosive behaviours and overfitting. The random gender effect offered no improvement over the fixed-effect specifications. Overall, M1 offered the best balance between parsimony and predictive accuracy.
5 Discussion
This study proposes a novel Bayesian hierarchical Beta regression model for analysing the spatio-temporal dynamics of obesity rates across Italian regions from 2010 to 2022. The results highlight pronounced gender differences, regional heterogeneity and dependence, and clear temporal trends, suggesting the importance of spatial, temporal, and demographic factors in explaining obesity patterns.
The choice of the Beta distribution is motivated not only by its support on the unit interval - making it particularly suitable for modelling rates and proportions - but also by its flexibility, as it accommodates a wide range of distributional shapes, depending on the parameter values (Johnson et al.,, 1995). In the present case study, the proposed Bayesian hierarchical Beta regression model incorporates structured spatial and temporal random effects additively, along with gender-specific intercepts. This model proves effective in capturing regional variations and temporal dependencies in obesity rates.
Some methodological limitations warrant consideration. The ICAR prior for the structured spatial random effects assumes complete spatial dependence, which may not fully align with the underlying spatial processes. Additionally, the additive specification of spatial and temporal random effects with separate priors may overlook potential space-time interactions. The exclusion of unstructured spatio-temporal variability, which could enhance flexibility in capturing localized deviations, represents another limitation. Furthermore, the regional-level analysis improves measurement reliability and mitigates data gaps but may obscure finer-scale variations at the provincial level. Despite these constraints, the study tries to advance the robustness and interpretability of spatio-temporal models, prioritizing methodological clarity and practical applicability. It contributes to the development of statistical tools for spatio-temporal analysis in public health and social sciences, offering a comprehensive understanding of obesity trends across Italian regions.
Our findings highlight the substantial role of the gender-specific random effects in estimating obesity outcomes, underscoring gender as a critical determinant of regional obesity rates. Spatial random effects reveal homogeneous patterns across the North-West, with the exception of Emilia-Romagna, which exhibits higher values. In the North-East, Friuli-Venezia Giulia and Veneto show similar values, while Trentino-Alto Adige displays lower magnitudes with a greater range. The Centre geographical zone is characterized by marked heterogeneity, with Lazio showing significantly lower values and Umbria higher ones. Southern regions generally exhibit higher spatial random effects, except for Calabria. The Islands show distinct spatial structures, an aspect enforced by the fact that Sicilia is treated as a neighbour to Calabria in the ICAR prior, while Sardegna is considered spatially isolated. Temporal random effects indicate a stable trend in the early years, followed by an increase in later years, suggesting a growing upward trend in obesity rates. Fixed effects, represented by exogenous covariates, appear to play a minimal role. Indeed, the application of the SSVS prior identifies only few covariates with significant associations to obesity, namely life expectancy, life expectancy in good health and overweight.
While this study does not establish causal relationships, it underscores the importance of modelling obesity trends through demographic and spatio-temporal dynamics rather than relying solely on fixed socioeconomic or lifestyle predictors. The findings reinforce the necessity of incorporating spatial and temporal structures in public health research. Future research could refine this approach by analysing finer spatial resolutions, such as provinces or municipalities, where fixed effects may be more relevant. Incorporating individual-level data and environmental factors could further enhance understanding of obesity dynamics, while non-parametric methods could facilitate data-driven regional clustering. Additionally, exploring interactions between spatial, temporal, and other random effects could provide deeper insights into regional health patterns.
In conclusion, this study demonstrates the effectiveness of spatio-temporal Bayesian modelling in capturing the complexities of obesity rates and in informing societal challenges using administrative data. It advocates for continued investigation of demographic, geographic, temporal, and behavioural factors to inform public health interventions and policy-making.
6 Competing interests
The authors declare that they have no competing interests.
7 Author contributions statement
L.R.: data collection and engineering, methodology, formal analysis, writing - original draft. R.A.: methodology, supervision, review & editing. M.C.: methodology, supervision, review & editing.
8 Data and code
The assembled dataset and the code used for implementing the analysis and running the models is available at the following GitHub repository:
https://github.com/LucianoRota00/Spatio-Temporal-Dynamics-of-Obesity-in-Italian-Regions.git.
9 Acknowledgments
L.R. gratefully acknowledges the financial support provided by the University of Bergamo for the Ph.D. programme in Statistics at the University of Milano-Bicocca.
{appendices}
10 Appendix A: description of the covariates
Table 7: Description of the covariates used in the analysis, grouped into four thematic categories: Socio-Demography, Mortality, Lifestyle, and General Wealth-related Information. The table reports the unit of measurement, availability by gender, and data source for each variable.
Variable Unit of measure By gender Source
Group 1 – Socio-Demography
Population count yes Eurostat
More than 65 years old percentage yes HFA
Foreigners percentage yes HFA
Foreigners from EU percentage yes HFA
More than 65 y/o living alone percentage yes HFA
Mean family components average no HFA
Unemployment percentage yes Eurostat
Gross Income euro (thousands) no Eurostat
Wedding rates percentage no HFA
University (from Bachelor’s to PhD) percentage yes Eurostat
Group 2 – Mortality
Cancer of digestive system percentage yes HFA
Cancer of stomach percentage yes HFA
Diabetes percentage yes HFA
Mental disorders percentage yes HFA
Blood diseases percentage yes HFA
Heart diseases percentage yes HFA
Digestive diseases percentage yes HFA
Liver diseases percentage yes HFA
Suicide percentage yes HFA
Group 3 – Lifestyle
Overweight percentage yes HFA
Overweight of younger percentage yes HFA
Cigarette consumption count yes HFA
Smoking rate percentage yes BES
Complete breakfast percentage yes HFA
Daily cheese consumption percentage yes HFA
Red meat consumption percentage yes HFA
Fish consumption percentage yes HFA
Daily vegetables consumption percentage yes HFA
Dinner as principal meal percentage yes HFA
No sport participation percentage yes HFA
Adequate nutrition percentage yes BES
Alcohol consumption percentage yes BES
Group 4 – General Wealth-related Information
Bad wealth percentage yes HFA
Life expectancy (LE) years yes HFA
LE in good wealth years yes HFA
LE without limitations years yes HFA
Drug consumption percentage yes HFA
Life satisfaction percentage yes BES
Health expenditure per capita euro/population no MEF
[…]
{tablenotes}
Source: HFA = Health for All database; BES = Italian Equitable and Sustainable Well-being report; MEF = Italian Ministry of Economy and Finance; Eurostat = Statistical Office of the European Union.
11 Appendix B: posterior summary statistics
This appendix provides the posterior summary statistics not reported in the main text.
Table 8: Posterior summary statistics for the region-specific spatial parameters $\psi_{i}$, with $i=1,\ldots,20$.
. Parameter (Region) Mean SD CV Skewness Kurtosis $\psi_{1}$ (Piemonte) -0.61 0.07 -0.12 0.18 3.46 $\psi_{2}$ (Valle d’Aosta) -0.48 0.09 -0.18 0.34 3.96 $\psi_{3}$ (Lombardia) -0.41 0.08 -0.19 0.25 3.41 $\psi_{4}$ (Liguria) -0.62 0.08 -0.12 0.11 3.42 $\psi_{5}$ (Emilia-Romagna) 0.25 0.08 0.33 0.32 3.59 $\psi_{6}$ (Trentino-Alto Adige) -0.49 0.12 -0.25 0.02 3.19 $\psi_{7}$ (Veneto) -0.02 0.07 -3.34 0.06 3.13 $\psi_{8}$ (Friuli-Venezia Giulia) 0.13 0.07 0.55 0.03 3.08 $\psi_{9}$ (Toscana) -0.15 0.08 -0.56 0.18 3.30 $\psi_{10}$ (Umbria) 0.29 0.07 0.26 0.19 3.29 $\psi_{11}$ (Marche) 0.03 0.07 2.73 0.03 3.09 $\psi_{12}$ (Lazio) -0.40 0.07 -0.18 0.20 3.55 $\psi_{13}$ (Abruzzo) 0.49 0.06 0.13 -0.03 3.05 $\psi_{14}$ (Molise) 0.89 0.08 0.09 -0.25 3.57 $\psi_{15}$ (Campania) 0.35 0.11 0.32 -0.31 3.65 $\psi_{16}$ (Puglia) 0.65 0.08 0.12 -0.55 4.34 $\psi_{17}$ (Basilicata) 0.70 0.09 0.13 -0.26 3.58 $\psi_{18}$ (Calabria) -0.07 0.12 -1.69 -0.13 3.12 $\psi_{19}$ (Sicilia) 0.16 0.10 0.60 -0.22 3.60 $\psi_{20}$ (Sardegna) -0.71 0.09 -0.12 -0.32 3.79
Table 9: Posterior summary statistics for time-specific parameters $\alpha_{t}$, with $t=1,\ldots,13$.
Parameter (Year) Mean SD CV Skewness Kurtosis
$\alpha_{1}$ (2010) -0.07 0.12 -1.83 -0.02 3.79
$\alpha_{2}$ (2011) -0.08 0.13 -1.65 0.11 3.64
$\alpha_{3}$ (2012) -0.08 0.14 -1.63 0.29 3.54
$\alpha_{4}$ (2013) -0.04 0.14 -3.51 0.26 3.36
$\alpha_{5}$ (2014) 0.04 0.14 3.13 0.37 3.44
$\alpha_{6}$ (2015) -0.06 0.14 -2.24 0.01 3.31
$\alpha_{7}$ (2016) 0.10 0.14 1.38 0.36 3.31
$\alpha_{8}$ (2017) 0.11 0.14 1.36 0.23 3.09
$\alpha_{9}$ (2018) 0.23 0.14 0.62 0.32 3.25
$\alpha_{10}$ (2019) 0.28 0.15 0.52 0.26 3.14
$\alpha_{11}$ (2020) 0.40 0.15 0.37 0.36 3.41
$\alpha_{12}$ (2021) 0.41 0.14 0.35 0.40 3.48
$\alpha_{13}$ (2022) 0.30 0.14 0.48 0.31 3.27
Table 10: Posterior summary statistics for precisions of the spatial and temporal random effects, $\tau_{\psi}$ and $\tau_{\alpha}$, respectively.
Parameter Mean SD CV Skewness Kurtosis
$\tau_{\psi}$ (Space precision) 2.37 0.82 0.34 0.80 4.14
$\tau_{\alpha}$ (Time precision) 100.84 65.26 0.65 2.19 12.08
[…]
Focusing on the precision parameters for the spatial and temporal random effects ($\tau_{\psi}$ and $\tau_{\alpha}$, respectively), we observe substantial Bayesian updating, particularly for the temporal component. Beginning with conservative Gamma priors, the posterior distribution of the spatial precision $\tau_{\psi}$ exhibits only a moderate increase. This restrained growth can be partly attributed to the ICAR prior structure, in which the full-conditional precision is scaled by the number of neighboring regions. In contrast, the temporal precision $\tau_{\alpha}$, whose prior specification does not involve such a multiplicative adjustment, undergoes a more pronounced shift, with a posterior mean markedly higher than that of $\tau_{\psi}$. However, direct comparison between these two parameters should be approached with caution, as their prior formulations differ substantially, rendering such comparisons potentially misleading. Both posterior distributions are characterized by high skewness and kurtosis, suggesting strong positive asymmetry and heavy tails.
12 Appendix C: predictive performance comparison
In this appendix we evaluate the predictive performance of four model specifications using a temporal training-test split. Specifically, the training set comprises data from year 1 to 12, while the test set consists of observations from 2022 ($t=13$), across all regions. As described in Section 4, we compare four Bayesian Beta regression models which differ in two aspects: (i) the specification of the gender effect (fixed vs. random effect), and (ii) the prior distribution assigned to the auto-regressive coefficient $\rho$. The resulting four model variants (named M1, M2, M3 and M4) are summarized in Table 11, with M1 corresponding to our main model described in Section 3.
With respect to gender, the random effect specification models the gender-specific intercept as
$\xi_{s}=\beta_{0}+\gamma_{s},\quad s=1,2,$
where $\beta_{0}$ is the global intercept and $\gamma_{s}$ represents a random deviation for each gender. The random effects follow a classical hierarchical formulation (e.g., Hoff,, 2009; Gelman,, 2006):
$\gamma_{s}\mid\mu_{s},\sigma^{2}_{s}\sim\mathcal{N}(\mu_{s},\sigma^{2}_{s}),\quad\mu_{s}\sim\mathcal{N}(0,4),\quad\sigma^{2}_{s}\sim U(1,10).$
In contrast, the gende fixed effect specification is the one we describe in Section 3.2.
As for the auto-regressive coefficient $\rho$, we consider two prior formulations: a $\text{Beta}(1,1)$ prior (constraining $\rho$ to the interval $(0,1)$) as specified in Section 3.2, and a standard normal prior $\mathcal{N}(0,1)$, which places no support restrictions and permits values outside the unit interval.
Table 11: Summary of model variants.
Model Gender Effect Prior on $\rho$
M1 Fixed $Beta(1,1)$
M2 Fixed $\mathcal{N}(0,1)$
M3 Random $Beta(1,1)$
M4 Random $\mathcal{N}(0,1)$
[…]
To evaluate the predictive performance, we compute standard error-based metrics for the test set: root mean squared error (RMSE), mean absolute error (MAE), and the sum of the log predictive densities (SLPD). Additionally, we include a measure we refer to as the Bayesian $p$-value (BPV), defined for each observation as the proportion of posterior predictive samples exceeding the observed value. Values of BPV near 0.5 indicate that the observation lies near the centre of the predicted distribution, suggesting a good fit. In contrast, values close to 0 or 1 suggest under or over prediction, respectively. We compute a mean BPV (MBPV) across all the test observations as a summary.
Table 12: Predictive performance metrics for the four model variants, evaluated on the 2022 test data. RMSE (F) and RMSE (M) refer to female and male subgroups, respectively.
Model RMSE RMSE (F) RMSE (M) MAE MBPV SLPD
M1 0.010 0.010 0.010 0.008 0.537 43.72
M2 0.013 0.012 0.014 0.010 0.717 33.17
M3 0.010 0.009 0.010 0.008 0.571 44.62
M4 0.014 0.013 0.015 0.012 0.757 27.78
[…]
Table 12 reports the predictive performance metrics for each model, evaluated on the 2022 test set. The results suggest that the prior specification for $\rho$ has a greater impact on predictive performance than the treatment of the gender effect. Specifically, models using a Normal prior for $\rho$ (M2, M4) tend to overestimate 2022 obesity rates. This is likely because the unbounded support of the Normal distribution allows $\rho$ to take values greater than one, permitting overly persistent or explosive temporal dynamics and leading to overfitting of upward trends in the training data. In contrast, the Beta(1,1)(1,1) prior restricts $\rho$ to the $(0,1)$ interval, constraining the model to more realistic levels of temporal autocorrelation. Models M1 and M3, which use this prior, yield more stable predictions and improved out-of-sample accuracy. Between M1 and M3, predictive performances are nearly identical. We select M1 for reporting in the main paper, as we think modeling gender as a random effect with only two levels (male and female) is not necessary. A fixed effect specification is more parsimonious in this case.